if oxygen at 128 Kpa is allowed to expand at constant temperature until its pressure is 101.3 kpa how much larger will the volume become?
1
Expert's answer
2020-06-08T15:40:30-0400

p1V1=p2V2;

V2/V1=p1/p2=128/101.3=1.263

Answer: the volume will be 1.263 times larger.